Hampton Industries had $58,000 in cash at year-end 2015 and $28,000 in cash at year-end 2016. The firm invested in property, plant, and equipment totaling $290,000. Cash flow from financing activities totaled +$160,000. Round your answers to the nearest dollar, if necessary. If accruals increased by $10,000, receivables and inventories increased by $170,000, and depreciation and amortization totaled $75,000, what was the firm's net income?
Answer #1
Decrease in cash = 58000-28000 =(30000)
Net cashflows from Operating activities = Cashoutflows from IA +Cashflows from financing activities -Decrease in cash balance
(290000)+160000 - 30000 = 100000
Net cash flows from Operating activities: 100000
Less: Increase in accruals -10000
Add: Increase in receivables 170000
Less: Depreciation -75000
Net income for the year 185000